OpenStudy (anonymous):

1+7y=5x-2 /formula and function problem (i need help)

jhonyy9 (jhonyy9):

5x-7y=3 x=(7y+3)/5 or 7y=5x-3 y=(5x-3)/7

OpenStudy (anonymous):

thanks